According to a government study among adults in the 25- to 34-year age group, the mean amount spent per year on reading and entertainment is $1,925. Assume that the distribution of the amounts spent follows the normal distribution with a standard deviation of $610. (Round z-score computation to 2 decimal places and your final answers to 2 decimal places.)

a. What percent of the adults spend more than $2,725 per year on reading and entertainment?

Percent %

b. What percent spend between $2,725 and $3,125 per year on reading and entertainment?

Percent %

c. What percent spend less than $1,300 per year on reading and entertainment?

Percent %

Answer by stanbon(75887)

According to a government study among adults in the 25- to 34-year age group, the mean amount spent per year on reading and entertainment is $1,925. Assume that the distribution of the amounts spent follows the normal distribution with a standard deviation of $610. (Round z-score computation to 2 decimal places and your final answers to 2 decimal places.)
a. What percent of the adults spend more than $2,725 per year on reading and entertainment?
-------
z(2725) = (2725-1925)/610 = 1.3115
------
P(x > 2725) = P(z > 1.3115) = normalcdf(1.3115,100) = 0.0948
-------------------------------------
Percent % = 9.48%
-----------------------
===============================
b. What percent spend between $2,725 and $3,125 per year on reading and entertainment?
Using the TI-84 normalcdf function::
normalcdf(2725,3125,1925,610) = 0.0703
Percent % = 7.03%
---------------------------
c. What percent spend less than $1,300 per year on reading and entertainment?
Percent %
Find the z value of 1300; Find the Probability of z less then that
z value.
------------
Cheers,
Stan H.
-----------